For a certain type of computer the length of time between charges of the battery is normally distributive with a mean of 50 hours and a standard deviation of 15 hours. Marvin owns one of these computers and wants to know the probability that the length time will be between 50 and 70 hours
Answer by edjones(7569)

z[1]=(50-50)/15=0
z[2]=(70-50)/15=20/15=4/3
0.409=area under the normal curve between 0 and 4/3
.
Ed